# Spatial exponential decay of perturbations in optimal control of general evolution equations

### v3

@@ 1. Introduction
Only recently, an exponential decay of sensitivities with regard to perturbations in space was considered and proven, e.g., in [19] for finite-dimensional nonlinear optimization problems on graphs under second order sufficient conditions that are uniform in the network size. Further, such an exponential sensitivity was leveraged in [20] for efficient neural-network-based approximations of separable optimal value functions. For linear-quadratic optimal control of stationary (such as Poisson and Helmholtz equation) and parabolic PDEs, an exponential decay of perturbations in space was shown in [10] under stabilizability and detectability assumptions, linking the decay in optimal control to system-theoretic properties.
In this work, we extend the result of the previous work [10] to general evolution equations governed by semigroups of operators. This class includes a large number of highly relevant systems (e.g. wave, transport, beam equations), which do not exhibit high regularity or a parabolic structure. Furthermore, our analysis features a weaker stabilizability/detectability assumption in the sense that we allow for the stabilized solution to exhibit an overshoot behaviour. On the contrary, in [10], such an overshoot behaviour is not admissible. The essence of the stabilizability/detectability assumption is, that the constants in the exponential estimate of the stabilized semigroups are uniform in the size of the spatial domain.
As a second contribution, we thoroughly investigate this assumption for several hyperbolic equations on an one-dimensional domain, namely the transport equation with constant and space-dependent velocity, the continuity equation and the wave equation. For all of these examples, we propose an easily-verifiable novel necessary and sufficient condition on the control and observation domain to ensure domain-uniform stabilizability/detectability.
@@ 1.1. Nomenclature
Let $\mathbb{N}$ denote the natural numbers, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, $\mathbb{R}_{\geq 0}=[0,\infty)$ and $\mathbb{R}_{>0}=(0,\infty)$. The Lebesgue measure of a measurable set $\Omega$ denoted by $|\Omega|$ and $L^{2}(\Omega)$ denotes the space of square integrable Lebesgue measurable functions. For a Hilbert space $X$ and $T>0$, we denote by $L^{p}(0,T;X),p\in[1,\infty)$, the space of $X$-valued $p$-Bochner-integrable functions endowed with the norm $\left\|f\right\|_{L^{p}(0,T;X)}^{p}:=\int_{0}^{T}\left\|f(t)\right\|_{X}^{p}\mathrm{d}t$. Further, $L^{\infty}(0,T;X)$ stands for the space of $X$-valued Bochner-integrable, essentially bounded functions with norm $\|\cdot\|_{L^{\infty}(0,T;X)}=\text{esssup}_{t\in[0,T]}\|f(t)\|$. The space of $X$-valued continuous functions is denoted by $C(0,T;X)$ with norm $\|\cdot\|_{C(0,T;X)}:=\max_{t\in[0,T]}\|f(t)\|$.
@@ 2. Problem statement
Correspondingly, the constraint and the state may be eliminated by above variation of constants formula of semigroup theory. Thus, via standard methods cf. [22], one may conclude that ($\mathrm{OCP}_{\Omega}^{T}$) has a unique solution $x_{\Omega}^{T}$ and $u_{\Omega}^{T}$, see, e.g., [18, Theorem 2.24]. If clear from context, we will mostly omit the indices $T$ and $\Omega$ for readability.
@@ 3. Exponential decay of optimal solutions in space
with $c\in L^{\infty}(\Omega_{L})$ essentially bounded from below and above by positive constants. Correspondingly, we may compute by means of the chain rule for weak derivatives (cf. also [10, Lemma 3]),
@@
As a second example, we consider distributed control of a wave equation on $\Omega\subset\mathbb{R}^{d}$, $d\in\{2,3\}$, given by
@@
e.g. on $L^{2}(\Omega)\times H^{-1}(\Omega)$ with $D(A_{\Omega,2})=H^{1}_{0}(\Omega)\times L^{2}(\Omega)$. Proceeding analogously to Example 7, however, yields a perturbation $S_{\Omega,1}$ that includes a first derivative, i.e., that is not bounded on $L^{2}(\Omega)$.
@@
on the state space $L^{2}(\Omega)\times L^{2}(\Omega;\mathbb{R}^{d})$ defined on a suitable subset of $H(\operatorname{div},\Omega)\times H^{1}(\Omega)$, in which the Dirichlet boundary condition on the displacement may be encoded by means of a unique reconstruction, cf. [17, Section 7.2]. Here, $H(\operatorname{div},\Omega)$ denotes the space of vector-valued functions for which the weak divergence exists in $L^{2}(\Omega)$. Importantly, this reformulation via $A_{\Omega,2}$ only includes first-order differential operators such that an analogous computation to Example 7 yields that $A_{\Omega,2}$ satisfies the assertions (i) and (iii) of Assumption 6.
@@
If $A_{\Omega}$ is given by a differential operator of order $r\geq 2$ the perturbation operators $S_{\Omega,i}^{\mu}$ are not bounded in $X_{\Omega}$ anymore, as they also involve derivative operators. Consequently, 6 is not satisfied in this case. In the previous Example 8 we were able to work around this issue via a change of variables. Nonetheless the natural question arises, if it is possible to relax 6 to the case of unbounded perturbation operators $S_{\Omega,i}^{\mu}:D(S_{\Omega,i}^{\mu})\rightarrow X_{\Omega}$. However, this leads to some heavy technical obstacles, the most important among which may be that the operators $A_{\Omega}+S_{1,\Omega}^{\mu}$ respectively $(A_{\Omega}+S_{2,\Omega}^{\mu})^{*}$ are not necessarily generators of strongly continuous semigroups anymore, see e.g. [4]. This generator property is, however, strictly necessary for the proof of our main result in order to ensure, that the disturbed optimality system is well-posed when written in scaled variables $\tilde{x}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}x(t)$, $\tilde{u}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}u(t)$, $\tilde{\lambda}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}\lambda(t)$. The question of what kind of perturbation operators $S_{\Omega}$ the $A_{\Omega}+S_{\Omega}$ remains a generator of a strongly continuous semigroup has been extensively studied in the literature. For bounded perturbations $S\in L(X_{\Omega})$ this is always true (see [6, Theorem 1.3]). For unbounded perturbation operators, there exists a variety of approaches: If $S_{\Omega}$ is bounded on $(D(A_{\Omega}),\left\|\cdot\right\|_{1})$ where $\left\|x\right\|_{1}:=\left\|(sI-A_{\Omega})\right\|{X_{\Omega}}$ for some $s\in\sigma(A)$ then $A_{\Omega}+S_{\Omega}$ is also an infinitesimal generator [6, Corollary 1.5]. If $S_{\Omega}$ is relatively $A_{\Omega}$-bounded, i.e.
$D(A_{\Omega})\subset D(S_{\Omega})\quad\textrm{and}\quad\exists\,a,b>0\,\forall x\in D(A_{\Omega}):\left\|S_{\Omega}x\right\|_{X_{\Omega}}\leq a\left\|Ax\right\|_{X_{\Omega}}+b\left\|x\right\|_{X_{\Omega}}$
@@
Having formulated suitable assumptions on the generator in Assumption 6 and on the control and observation operator in Assumption 4 of the optimal control problem ($\mathrm{OCP}_{\Omega}^{T}$), we may now state our main result of this part. Therein, we prove the spatial decay of perturbations of the optimality system (1). In its formulation, we assume a uniform bound on the solution operator that will be verified under domain-uniform stabilizability and detectability assumptions in the subsequent Theorem 13. The formulation and the proof of both results is similar to the results considering exponential decay in time [12] or spatial decay in elliptic and parabolic equations [10].
@@
both with Dirichlet boundary conditions $(x_{1}(t))(0)=(x_{1}(t))(L)=0$. Let $x_{L}^{i},i\in\{1,2\}$ be classical (in particular continuous in space) solutions of these two formulations. Assume domain-uniform exponential decay of the state of this equation around some point in space, i.e., there exist $P\in\mathbb{R}$ and $M,\mu>0$ such that
@@
The proof follows along the lines of [12, Theorem 10] and is provided for completeness in Appendix A. ∎
@@ 4. Domain-uniform stabilizability of the transport equation on a scalar domain
where the measurement domain $\Omega_{o}\subset\mathbb{R}_{\geq 0}$ is assumed to be Lebesgue-measurable with positive measure and for $L>0$ we define $\Omega_{L}^{o}:=\Omega_{o}\cap[0,L]$. For $\Omega_{L}^{o}=\Omega_{L}^{c}$ we have $B_{\Omega_{L}}^{*}=C_{\Omega_{L}}$. The theorem of Stone [24, Theorem 3.8.6] implies that the operator $A_{\Omega_{L}}$ generates a unitary group since it is skew-adjoint. Note that due to duality domain-uniform detectability of $(A,C)$ is the same as domain-uniform stabilizability of $(A^{*},C^{*})$.
@@ 4.2. Characterization of control domains for domain-uniform stabilizability
19 (ii) provides a simple algebraic characterization of control domains which ensure domain-uniform stabilizability of (11). This characterization has two main advantages: First, it is a useful tool in order to check domain-uniform stabilizability for a given control domain. Second, it even allows for the iterative construction of such a control domain. For example, this can be done by predefining constants $K>k>0$ and - starting from $a_{1}=0$ - successively choosing $a_{n}$ and $b_{n}$ such that condition (ii) is fulfilled for all $m\leq n$. In (iii) this characterization is rephrased to allow for an interpretation in terms of measures. This characterization shows that the size of the control domain $|\Omega_{c}|$ has to grow at least linearly with the domain size $|\Omega|$ in order to guarantee domain-uniform stabilizability. In [10, Eq.(3.18)] a similar condition is used for this purpose. However in this work the authors assume that the uncontrollable domain, i.e. $\Omega_{L}\setminus\Omega_{L}^{c}$, has $L$-uniformly bounded Lebesgue measure. Here, we stress that condition (iii) in 19 actually allows for the set $\mathbb{R}_{\geq}\setminus\Omega_{c}$ to be unbounded as long as the Lebesgue measure $|\Omega_{L}\setminus\Omega_{L}^{c}|$ does not grow more than linearly with the domain size $L$. In this sense 19(iii) is more general than the previous result of [10].
@@
Therefore the operator $A_{\Omega_{L}}:\mathrm{dom}(A_{\Omega_{L}})\rightarrow L^{2}([0,L])$ is a formal adjoint operator of $A_{\Omega_{L}}$. Since $\mathrm{dom}(A_{\Omega_{L}})$ is a dense subset of $L^{2}([0,L])$ it can be extended to a maximal formal adjoint operator which is the adjoint operator [24, Section 2.8]. This shows $\mathrm{dom}(A_{\Omega_{L}})\subset\mathrm{dom}(A_{\Omega_{L}}^{*})\subset H_{1}([0,L])$. Let $v\in H_{1}([0,L])$ such that for all $x\in\mathrm{dom}(A_{\Omega_{L}})$
@@ 6.1. System formulation and unitary group
with time horizon $T>0$, control $u\in L^{2}(\Omega_{L}^{c}\times[0,T])=:U_{\Omega_{L}}$, initial distributions $x_{\Omega_{L}}^{0}\in H^{2}(\Omega_{L})\cap H_{0}^{1}(\Omega_{L})$ and $x_{\Omega_{L}}^{1}\in H^{1}(\Omega_{L})$ and velocity $c>0$. We first consider the uncontrolled equation ($u\equiv 0$). In this case, the wave equation can be rewritten as a first-order PDE of the form $\dot{z}=Az$ where $z=\begin{pmatrix}x&\frac{\partial}{\partial t}x\end{pmatrix}^{\top}$. For this purpose, we define the Dirichlet Laplacian via
@@
From [24, Proposition 3.6.1] we know that $A_{\Omega_{L}}^{0}$ is strictly positive, i.e. $A_{\Omega_{L}}^{0}=(A_{\Omega_{L}}^{0})^{*}$ and
@@
[24, Proposition 3.7.6] tells us, that $A$ is skew-adjoint. Therefore, we can apply the Theorem of Stone [24, Prop. 3.8.6] to show, that $A$ generates a unitary group $\left(T_{L}(t)\right)_{t\in\mathbb{R}}$ on $H_{0}^{1}(\Omega_{L})\times L^{2}(\Omega_{L})$. It is easy to see that $x\in D(A_{\Omega_{L}}^{0})$ is a classical solution of the uncontrolled wave equation if and only if $z=\begin{pmatrix}x&\frac{\partial}{\partial t}x\end{pmatrix}^{\top}\in D(A)$ is a classical solution of
@@
Since in this case $z(t)=T_{L}(t)z_{0}$ (see [6, Proposition 6.2]) and $\left(T_{L}(t)\right)_{t\in\mathbb{R}}$ is a unitary group we find the equality (see also [24, Proposition 3.8.7])

### v4

@@ 1. Introduction
Only recently, an exponential decay of sensitivities with regard to perturbations in space was considered and proven, e.g., in [18] for finite-dimensional nonlinear optimization problems on graphs under second order sufficient conditions that are uniform in the network size. Further, such an exponential sensitivity was leveraged in [19] for efficient neural-network-based approximations of separable optimal value functions. For linear-quadratic optimal control of stationary (such as Poisson and Helmholtz equation) and parabolic PDEs, an exponential decay of perturbations in space was shown in [Göttlich2024] under stabilizability and detectability assumptions, linking the decay in optimal control to system-theoretic properties.
In this work, we extend the result of the previous work [Göttlich2024] to general evolution equations governed by semigroups of operators. This class includes a large number of highly relevant systems (e.g. wave, transport, beam equations), which do not exhibit high regularity or a parabolic structure. Furthermore, our analysis features a weaker stabilizability/detectability assumption in the sense that we allow for the stabilized solution to exhibit an overshoot behaviour. On the contrary, in [Göttlich2024], such an overshoot behaviour is not admissible. The essence of the stabilizability/detectability assumption is, that the constants in the exponential estimate of the stabilized semigroups are uniform in the size of the spatial domain.
As a second contribution, we thoroughly investigate this assumption for several hyperbolic equations on a one-dimensional domain, namely the transport equation with constant and space-dependent velocity, the continuity equation and the wave equation. For all of these examples, we propose an easily-verifiable novel necessary and sufficient condition on the control and observation domain to ensure domain-uniform stabilizability/detectability.
@@ 1.1. Nomenclature
Let $\mathbb{N}$ denote the natural numbers, $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$, $\mathbb{R}_{\geq 0}=[0,\infty)$ and $\mathbb{R}_{>0}=(0,\infty)$. The Lebesgue measure of a measurable set $\Omega$ denoted by $|\Omega|$ and $L^{2}(\Omega)$ denotes the space of square integrable Lebesgue measurable functions. For a Hilbert space $X$ and $T>0$, we denote by $L^{p}(0,T;X),p\in[1,\infty)$, the space of $X$-valued $p$-Bochner-integrable functions endowed with the norm $\left\|f\right\|_{L^{p}(0,T;X)}^{p}:=\int_{0}^{T}\left\|f(t)\right\|_{X}^{p}\mathrm{d}t$. Further, $L^{\infty}(0,T;X)$ stands for the space of $X$-valued Bochner-integrable, essentially bounded functions with norm $\|\cdot\|_{L^{\infty}(0,T;X)}=\text{esssup}_{t\in[0,T]}\|f(t)\|$. The space of $X$-valued continuous functions over $[0,T]$ is denoted by $C(0,T;X)$ with norm $\|\cdot\|_{C(0,T;X)}:=\max_{t\in[0,T]}\|f(t)\|$.
@@ 2. Problem statement
Correspondingly, the constraint and the state may be eliminated by above variation of constants formula of semigroup theory. Thus, via standard methods cf. [Tröltzsch2010], one may conclude that ($\mathrm{OCP}_{\Omega}^{T}$) has a unique solution $x_{\Omega}^{T}$ and $u_{\Omega}^{T}$, see, e.g., [17, Theorem 2.24]. If clear from context, we will mostly omit the indices $T$ and $\Omega$ for readability.
@@ 3. Exponential decay of optimal solutions in space
with $c\in L^{\infty}(\Omega_{L})$ essentially bounded from below and above by positive constants. Correspondingly, we may compute by means of the chain rule for weak derivatives (cf. also [Göttlich2024, Lemma 3]),
@@
As a second example, we consider distributed control of a wave equation on $\Omega\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, given by
@@
e.g. on $L^{2}(\Omega)\times H^{-1}(\Omega)$ with $D(A_{\Omega,1})=H^{1}_{0}(\Omega)\times L^{2}(\Omega)$. Proceeding analogously to Example 7, however, yields a perturbation $S_{\Omega,1}$ that includes a first derivative, i.e., that is not bounded on $L^{2}(\Omega)$.
@@
on the state space $L^{2}(\Omega)\times L^{2}(\Omega;\mathbb{R}^{d})$ defined on a suitable subset of $H(\operatorname{div},\Omega)\times H^{1}(\Omega)$, in which the Dirichlet boundary condition on the displacement may be encoded by means of a unique reconstruction, cf. [16, Section 7.2]. Here, $H(\operatorname{div},\Omega)$ denotes the space of vector-valued functions for which the weak divergence exists in $L^{2}(\Omega)$. Importantly, this reformulation via $A_{\Omega,2}$ only includes first-order differential operators such that an analogous computation to Example 7 yields that $A_{\Omega,2}$ satisfies the assertions (i) and (iii) of Assumption 6.
@@
If $A_{\Omega}$ is given by a differential operator of order $r\geq 2$ the perturbation operators $S_{\Omega,i}^{\mu}$ are not bounded in $X_{\Omega}$ anymore, as they also involve derivative operators. Consequently, 6 is not satisfied in this case. In the previous Example 8 we were able to work around this issue via a change of variables. Nonetheless the natural question arises, if it is possible to relax 6 to the case of unbounded perturbation operators $S_{\Omega,i}^{\mu}:D(S_{\Omega,i}^{\mu})\rightarrow X_{\Omega}$. However, this leads to some heavy technical obstacles, the most important among which may be that the operators $A_{\Omega}+S_{1,\Omega}^{\mu}$ respectively $(A_{\Omega}+S_{2,\Omega}^{\mu})^{*}$ are not necessarily generators of strongly continuous semigroups anymore, see e.g. [4]. This generator property is, however, strictly necessary for the proof of our main result in order to ensure, that the disturbed optimality system is well-posed when written in scaled variables $\tilde{x}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}x(t)$, $\tilde{u}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}u(t)$, $\tilde{\lambda}(t):=e^{\mu\left\|P-\cdot\right\|_{1}}\lambda(t)$. The question of what kind of perturbation operators $S_{\Omega}$ the $A_{\Omega}+S_{\Omega}$ remains a generator of a strongly continuous semigroup has been extensively studied in the literature. For bounded perturbations $S\in L(X_{\Omega})$ this is always true (see [6, Theorem 1.3]). For unbounded perturbation operators, there exists a variety of approaches: If $S_{\Omega}$ is bounded on $(D(A_{\Omega}),\left\|\cdot\right\|_{1})$ where $\left\|x\right\|_{1}:=\left\|(sI-A_{\Omega})\right\|{X_{\Omega}}$ for some $s\in\rho(A)$ then $A_{\Omega}+S_{\Omega}$ is also an infinitesimal generator [6, Corollary 1.5]. If $S_{\Omega}$ is relatively $A_{\Omega}$-bounded, i.e.
$D(A_{\Omega})\subset D(S_{\Omega})\quad\textrm{and}\quad\exists\,a,b>0\,\forall x\in D(A_{\Omega}):\left\|S_{\Omega}x\right\|_{X_{\Omega}}\leq a\left\|A_{\Omega}x\right\|_{X_{\Omega}}+b\left\|x\right\|_{X_{\Omega}}$
@@
Having formulated suitable assumptions on the generator in Assumption 6 and on the control and observation operator in Assumption 4 of the optimal control problem ($\mathrm{OCP}_{\Omega}^{T}$), we may now state our main result of this part. Therein, we prove the spatial decay of perturbations of the optimality system (1). In its formulation, we assume a uniform bound on the solution operator that will be verified under domain-uniform stabilizability and detectability assumptions in the subsequent Theorem 13. The formulation and the proof of both results is similar to the results considering exponential decay in time [11] or spatial decay in elliptic and parabolic equations [Göttlich2024].
@@
both with Dirichlet boundary conditions $(x_{1}(t))(0)=(x_{1}(t))(L)=0$. Let $x_{L}^{i},i\in\{1,2\}$ be classical (in particular continuous in space) solutions of these two formulations. We write $x_{L}^{i}(\omega,t)$ instead of $(x_{L}^{i}(t))(\omega)$ where $\omega$ is the spatial variable of the solution. Assume domain-uniform exponential decay of the state of this equation around some point in space, i.e., there exist $P\in\mathbb{R}$ and $M,\mu>0$ such that
@@
The proof follows along the lines of [11, Theorem 10] and is provided for completeness in Appendix A. ∎
@@ 4. Domain-uniform stabilizability of the transport equation on a scalar domain
where the measurement domain $\Omega_{o}\subset\mathbb{R}_{\geq 0}$ is assumed to be Lebesgue-measurable with positive measure and for $L>0$ we define $\Omega_{L}^{o}:=\Omega_{o}\cap[0,L]$. For $\Omega_{L}^{o}=\Omega_{L}^{c}$ we have $B_{\Omega_{L}}^{*}=C_{\Omega_{L}}$. The theorem of Stone [22, Theorem 3.8.6] implies that the operator $A_{\Omega_{L}}$ generates a unitary group since it is skew-adjoint. Note that due to duality domain-uniform detectability of $(A,C)$ is the same as domain-uniform stabilizability of $(A^{*},C^{*})$.
@@ 4.2. Characterization of control domains for domain-uniform stabilizability
19 (ii) provides a simple algebraic characterization of control domains which ensure domain-uniform stabilizability of (11). This characterization has two main advantages: First, it is a useful tool in order to check domain-uniform stabilizability for a given control domain. Second, it even allows for the iterative construction of such a control domain. For example, this can be done by predefining constants $K>k>0$ and - starting from $a_{1}=0$ - successively choosing $a_{n}$ and $b_{n}$ such that condition (ii) is fulfilled for all $m\leq n$. In (iii) this characterization is rephrased to allow for an interpretation in terms of measures. This characterization shows that the size of the control domain $|\Omega_{c}|$ has to grow at least linearly with the domain size $|\Omega|$ in order to guarantee domain-uniform stabilizability. In [Göttlich2024, Eq.(3.18)] a similar condition is used for this purpose. However in this work the authors assume that the uncontrollable domain, i.e. $\Omega_{L}\setminus\Omega_{L}^{c}$, has $L$-uniformly bounded Lebesgue measure. Here, we stress that condition (iii) in 19 actually allows for the set $\mathbb{R}_{\geq}\setminus\Omega_{c}$ to be unbounded as long as the Lebesgue measure $|\Omega_{L}\setminus\Omega_{L}^{c}|$ does not grow more than linearly with the domain size $L$. In this sense 19(iii) is more general than the previous result of [Göttlich2024].
@@
Therefore the operator $A_{\Omega_{L}}:\mathrm{dom}(A_{\Omega_{L}})\rightarrow L^{2}([0,L])$ is a formal adjoint operator of $A_{\Omega_{L}}$. Since $\mathrm{dom}(A_{\Omega_{L}})$ is a dense subset of $L^{2}([0,L])$ it can be extended to a maximal formal adjoint operator which is the adjoint operator [22, Section 2.8]. This shows $\mathrm{dom}(A_{\Omega_{L}})\subset\mathrm{dom}(A_{\Omega_{L}}^{*})\subset H_{1}([0,L])$. Let $v\in H_{1}([0,L])$ such that for all $x\in\mathrm{dom}(A_{\Omega_{L}})$
@@ 6.1. System formulation and unitary group
with time horizon $T>0$, control $u\in L^{2}(\Omega_{L}^{c}\times[0,T])=:U_{\Omega_{L}}$, initial distributions $x_{\Omega_{L}}^{0}\in H^{2}(\Omega_{L})\cap H_{0}^{1}(\Omega_{L})$ and $x_{\Omega_{L}}^{1}\in H^{1}(\Omega_{L})$ and velocity $c>0$. We first consider the uncontrolled equation ($u\equiv 0$). In this case, the wave equation can be rewritten as a first-order PDE of the form $\dot{z}=A_{\Omega_{L}}z$ where $z=\begin{pmatrix}x&\frac{\partial}{\partial t}x\end{pmatrix}^{\top}$. For this purpose, we define the Dirichlet Laplacian via
@@
From [22, Proposition 3.6.1] we know that $A_{\Omega_{L}}^{0}$ is strictly positive, i.e. $A_{\Omega_{L}}^{0}=(A_{\Omega_{L}}^{0})^{*}$ and
@@
[22, Proposition 3.7.6] tells us, that $A$ is skew-adjoint. Therefore, we can apply the Theorem of Stone [22, Prop. 3.8.6] to show, that $A$ generates a unitary group $\left(T_{L}(t)\right)_{t\in\mathbb{R}}$ on $H_{0}^{1}(\Omega_{L})\times L^{2}(\Omega_{L})$. It is easy to see that $x\in D(A_{\Omega_{L}}^{0})$ is a classical solution of the uncontrolled wave equation if and only if $z=\begin{pmatrix}x&\frac{\partial}{\partial t}x\end{pmatrix}^{\top}\in D(A)$ is a classical solution of
@@
Since in this case $z(t)=T_{L}(t)z_{0}$ (see [6, Proposition 6.2]) and $\left(T_{L}(t)\right)_{t\in\mathbb{R}}$ is a unitary group we find the equality (see also [22, Proposition 3.8.7])